OpenStudy (anonymous):

Is the inequality sometimes, always, or never true -2(2x + 9)> -4x + 9 2. Solve the compound inequality. Graph the solution. -2 <= 2x - 4 < 4 0 <= x < -2 1 <= x < 4 1 <= x < 0 3 <= x < 6

OpenStudy (jdoe0001):

so solve -2(2x + 9)> -4x + 9 and find out :)

OpenStudy (cwrw238):

-2(2x + 9)> -4x + 9 -4x - 18 > -4x + 9 what do you think?

OpenStudy (anonymous):

sometimes true idk honestly because this is my weakest part is inequalities

OpenStudy (jdoe0001):

why not just solve it?

OpenStudy (jdoe0001):

how would you solve say => -4x - 18 = -4x + 9

OpenStudy (cwrw238):

add 4x to both sides of the inequality - what do you get?

OpenStudy (anonymous):

-8x

OpenStudy (cwrw238):

the rules for inequalities are similar to those for equations

OpenStudy (cwrw238):

-4x + 4x = -8x????

OpenStudy (cwrw238):

nope

OpenStudy (anonymous):

Oh no sorry x

OpenStudy (cwrw238):

-1 + 1 = 0 -2 + 2 = 0 -x + x = 0

OpenStudy (cwrw238):

-4x + 4x = ?

OpenStudy (anonymous):

so it is always equal

OpenStudy (cwrw238):

it should be obvious - compare - 4x + 4x = 0

OpenStudy (cwrw238):

no go back to the inequality -4x - 18 > -4x + 9 after adding 4x to both sides we get -18 > + 9 is that correct??

OpenStudy (anonymous):

yes

OpenStudy (cwrw238):

is -18 greater than + 9 ?

OpenStudy (anonymous):

no

OpenStudy (cwrw238):

right so whats the answer to the original question sometimes, never or always?

OpenStudy (anonymous):

never

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

ok